Diary

20150820_23xxxx PockelsCell145 installed in BlueChannel replacing PockelsCell144
20150821_150055 PCT measurement
20150821_163526 PCD off
20150821_174238 Low flux regime, loops on. Snapshot
20150821_174600 Chiller halted. Automatic poweroff
20150821_18xxxx Test of MoCon ethernet port
20150821_184600 Rack powercycle. Power up
20150821_192318 Red pupil ear. Snapshot
20150821_192429 PCD on
20150821_192547 PCT measurement
20150821_195819 Pupil stabilization loop test
20150821_215603 PCD off
20150821_224723 PCD on
20150821_224924 Pupil stabilization loop test
20150822_021200 Device Error Pupil Motor Red
20150822_022500 Patrol Camera disconnected and re-plugged after about 30min. After that device error on MoCon: SW restarted
20150822_170222 PCT measurement - Red Pupil Ear disappeared
20150822_202506 PCT measurement. Ok
20150822_202800 shutdown-restart-powerup
20150822_205600 SW restart
20150823_165520 PCT measurement. Ok
20150823_170000 Poweroff. We need the electronics for dx-lgsw

Zabbix

Zabbix patched to allow for 5min zoom in graphs. Downloaded a patch called zabbix-2.0.7-zoom.path from the web and installed on both WorkStation1 and WorkStation2

Pupil stabilization loop stability with flux variations

Between 20150821_195819 and 224924 we run with pupil and jitter loop closed and different values of flux to see if the algorithms have a dependence on the flux. The short answer is no.

The flux is adjusted, we wait for the loops to converge and we store pupil motor position and piezo mirror offset. We repeat for several flux levels. We never saw an abrupt change of pupil motor position or piezo mirror offset after the flux was adjusted, and this is considered an argument to assert that the pupil stabilization algorithm is robust to flux variations.

We could instead notice a slow drift with 1h timescale. The PtV drift over a 2h range are the following:

Ch (x,y) Pupil Motor (steps) Piezo (mrad)
Blue 5, 7 0.2, 0.2
Yellow 11, 12 0.5, 0.3
Red 18, 17 0.5, 0.7

Pockels cells

System has been long all night long without flaws.

Pockels cells are still ok: they have been on continuosly from 20150821_000519 to 20150821_163526 and no issues in the PCT have been seen. We tested both normal humidity (30%) and low humidity regimes.

At 20150821_163526 we powered off the pockel cell driver to check the pupil stabilization routine in low flux regime

At 192000 after a complete power-cycle of the system the red pupil has "ears"

The system is left on continously most of the time with PCD on and dry air. At 20150822_170222 the "ears" are disappeared and don't reappear until the end of the test on 20150823_170000

Topic revision: r7 - 27 Aug 2015, LorenzoBusoni
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback